solve dx/(y-z)=dy/(z-x)=dz/(x-y)



1
Expert's answer
2021-06-02T17:05:13-0400

"(1dx+mdy+ndz)\/(1P+mQ+nR)=0"

"1,m,n, are \, multipliers"

"1dx+1dy+1dz=0"

"1(y-z)+1(z-x)+1(x-y)=0"

"x(y-z) +y(z-y) +z(x-y) =0" "=\int (1dx+1dy+1dy)=0"

"x+y+z=C1"

"\int (xdx+ydy+zdz) =0"

"x^2+y^2+z^2=C2"